OpenStudy (anonymous):

find the slope and y intercept f(x)=-5x-4

OpenStudy (aroub):

slope=-5

OpenStudy (aroub):

y-intercept=-4

OpenStudy (saifoo.khan):

When x intercpts y = 0 When y intercpts x = 0

OpenStudy (saifoo.khan):

Slope is the number with x.

OpenStudy (saifoo.khan):

In this case, slope is -5. y-intercpt = -5(0)-4 -4

OpenStudy (anonymous):

y = mx +c is the slope intercept form, here m is the slope and b is the y intercept f(x) =-5x -4. Compare m =-5 and b = -4